# Metadata-enhanced contrastive learning from retinal optical coherence tomography images

### v2

@@ 3.3.4 Downstream tasks
    Dense retinal surface segmentation, where models simultaneously regress the column-wise location of eleven retinal cell layers along each OCT A-scan as in (He et al., 2021). Unlike the aforementioned image-level tasks, this task requires learning pixel-level structural variations which are known to be affected by pathology, healthy ageing and differences in patient sex.
@@ 4.1 Self-supervised pretraining on retinal OCT data is crucial for downstream performance
In linear evaluation, RETFound, which was trained as a Masked Autoencoder to reconstruct images on a patch-level, outperformed all other methods on retinal layer segmentation. This reflects other works suggesting that image-level representation learning tasks, such as contrastive learning, do not generalise well to dense pixel-level tasks like segmentation (Wang et al., 2021; Li et al., 2022; Hu et al., 2021; Chaitanya et al., 2020).
@@
Southampton Eye Unit
Moorfields Eye Hospital
Patient age regression Refer to caption Visual acuity regression Refer to caption Late vs. Early AMD classification Refer to caption Dry vs. Wet AMD classification Refer to caption
Patient sex classification Refer to caption Retinal layer segmentation Refer to caption Early AMD vs. Healthy classification Refer to caption
Patient age regression Refer to caption Visual acuity regression Refer to caption Late vs. Early AMD classification Refer to caption
Patient sex classification Refer to caption Retinal layer segmentation Refer to caption Refer to caption Refer to caption
@@ 5 Discussion
Our study represents the first to fully establish the wide ranging potential of metadata in self-supervised pretraining in retinal OCT. Previous works in medical domains (Chen et al., 2021; Azizi et al., 2021; Taleb et al., 2022; Zeng et al., 2021a; Vu et al., 2021) have either focussed on evaluating multiple downstream tasks or on varying labelled subset sizes. Our work uses seven tasks, two datasets and seven labelled subset sizes to provide a comprehensive picture of in-domain self-supervised pretraining. Furthermore while related works focused on improving a single contrastive loss, we used both SimCLR and BYOL to ensure that pretraining with metadata represents a general improvement to constrastive learning as a whole, rather than addressing the flaws of a single framework. Overall, our experiments more closely resemble those used in studies on natural images where more exhaustive evaluations are standard practise.
@@
To address some of these issues future work could, for each contrastive pair of images, compare the auto-generated metadata relationship reported by Equation 1 to a fully unsupervised similarity score computed using standard contrastive pretraining. Then, by highlighting cases where both BYOL and SimCLR disagree with Equation 1 we could begin to identify any systematic errors in our formulation.
@@ A.1 Unfrozen finetuning results
Southampton Eye Unit
Moorfields Eye Hospital
Patient age regression Refer to caption Visual acuity regression Refer to caption Late vs. Early AMD classification Refer to caption Dry vs. Wet AMD classification Refer to caption
Patient sex classification Refer to caption Retinal layer segmentation Refer to caption Early AMD vs. Healthy classification Refer to caption
Patient age regression Refer to caption Visual acuity regression Refer to caption Late vs. Early AMD classification Refer to caption
Patient sex classification Refer to caption Retinal layer segmentation Refer to caption Refer to caption Refer to caption
Figure 5: Performance of finetuning models with unfrozen weights on downstream tasks on both datasets. Depicted is the performance (with 95% CIs) against varying sizes of the labelled subsets used for finetuning.
@@ A.2 Contrastive pretraining details
In BYOL we choose the update coefficient $\tau=0.9995$ in accordance with the authors recommendations for smaller batch sizes. We also found that for metadata-enhanced SimCLR setting $Q=20$ using the configuration described in (Chuang et al., 2020) increased stability during pretraining.
@@ A.5 Robustness tests
Our metadata pairings defined in Equation 1 permit only positive relationships between images from the same longitude. This introduces a theoretical degenerate solution in which images from the same eye are mapped to the same latent feature point. In analogy to BYOL (Grill et al., 2020) we find that this degenerate solution reliably does not occur (see Figure 6). Moreover, as expected relationships between embeddings of images of the same eye over time respect clinical expectations. Namely, embeddings of images that were taken further apart in time, over which more progression can occur, are typically more distant. While this trend holds in most cases, it highlights two exceptions. Firstly, some images taken over a small window are in fact distant in feature space. This reflects that AMD can progress very rapidly in a short space of time, such as conversion to the late stage. Conversely, some patients progress very slowly and travel a relatively short distance in feature space even over a long time period.
@@
ZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ 3.3.4 Downstream tasks
    Dense retinal surface segmentation, where models simultaneously regress the column-wise location of eleven retinal cell layers along each OCT A-scan as in [23]. Unlike the aforementioned image-level tasks, this task requires learning pixel-level structural variations which are known to be affected by pathology, healthy ageing and differences in patient sex.
@@ 4.1 Self-supervised pretraining on retinal OCT data is crucial for downstream performance
In linear evaluation, RETFound, which was trained as a Masked Autoencoder to reconstruct images on a patch-level, outperformed all other methods on retinal layer segmentation. This reflects other works suggesting that image-level representation learning tasks, such as contrastive learning, do not generalise well to dense pixel-level tasks like segmentation [44, 31, 25, 6].
@@ 5 Discussion
Our study represents the first to fully establish the wide ranging potential of metadata in self-supervised pretraining in retinal OCT. Previous works in medical domains [11, 2, 42, 48, 43] have either focussed on evaluating multiple downstream tasks or on varying labelled subset sizes. Our work uses seven tasks, two datasets and seven labelled subset sizes to provide a comprehensive picture of in-domain self-supervised pretraining. Furthermore while related works focused on improving a single contrastive loss, we used both SimCLR and BYOL to ensure that pretraining with metadata represents a general improvement to constrastive learning as a whole, rather than addressing the flaws of a single framework. Overall, our experiments more closely resemble those used in studies on natural images where more exhaustive evaluations are standard practise.
@@
To address some of these issues future work could, for each contrastive pair of images, compare the auto-generated metadata relationship reported by Equation 1 to a fully unsupervised similarity score computed using standard contrastive pretraining. Then, by highlighting cases where both BYOL and SimCLR disagree with Equation 1 we could begin to identify any systematic errors in our formulation. Furthermore, future work can begin to investigate the potentially utility to prognostic predictions of disease progression through embedding temporal dynamics into pretraining [38].
@@ A.1 Unfrozen finetuning results
Figure 5: Performance of finetuning models with fully unfrozen weights on downstream tasks on both datasets. Depicted is the performance (with 95% CIs) against varying sizes of the labelled subsets used for finetuning.
@@ A.2 Contrastive pretraining details
In BYOL we choose the update coefficient $\tau=0.9995$ in accordance with the authors recommendations for smaller batch sizes. We also found that for metadata-enhanced SimCLR setting $Q=20$ using the configuration described in [Chuang et al., 2020] increased stability during pretraining.
@@ A.5 Robustness tests
Our metadata pairings defined in Equation 1 permit only positive relationships between images from the same longitude. This introduces a theoretical degenerate solution in which images from the same eye are mapped to the same latent feature point. In analogy to BYOL [Grill et al., 2020] we find that this degenerate solution reliably does not occur (see Figure 6). Moreover, as expected relationships between embeddings of images of the same eye over time respect clinical expectations. Namely, embeddings of images that were taken further apart in time, over which more progression can occur, are typically more distant. While this trend holds in most cases, it highlights two exceptions. Firstly, some images taken over a small window are in fact distant in feature space. This reflects that AMD can progress very rapidly in a short space of time, such as conversion to the late stage. Conversely, some patients progress very slowly and travel a relatively short distance in feature space even over a long time period.